Chloe has been offered positions by two car dealers. The first company pays a salary of $19,000 plus a commission of $500 for each car sold. The second pays a salary of $14,000 plus a commission of $700 for each car sold. How many cars would need to be sold to make the total pay the same?

Answer by Theo(13342)

equation for first company:
y = 19000 + 500 * x
equation for second company:
y = 14000 + 700 * x
y is the total compensation.
x is the number of cars sold.
the compensation is equal when the first equation is equal to the second.
you get:
19000 + 500 * x = 14000 + 700 * x
subtract 14000 from both sides and subtract 500 from both sides to get:
5000 = 200 * x
solve for x to get:
x = 5000 / 200 = 25
the compensation will be the same when 25 cars are sold.
19000 + 500 * 25 = 31500
14000 + 700 * 25 = 31500